# Does Diffusion Beat GAN in Image Super Resolution?

### v1

@@ Appendix A Qualitative Comparison
Appendix B Architecture details
@@
Appendix C Training details
We use a linear warmup schedule over 1000 training steps to train the Diffusion model on image crops, followed by a constant learning rate of 3e-5. For full-resolution inputs, we decrease the learning rate to 2e-6 and increase the duration of the warmup phase to 10k training steps.
@@ Appendix C Training details
140k iterations for pretraining and training with the adversarial loss for the GAN-based SR model take approximately three days. Training of Diffusion models takes $\sim 1$a months with the same resources.
@@
Appendix D Human evaluation
Our main tool for training early stopping is side-by-side (SbS) comparisons of Super Resolution results between two subsequent checkpoints of the same model. For that, we use an internal crowdsourcing platform with non-expert assessors. Before labeling, each candidate assessor must pass an exam and achieve at least an 80% After that, we ask assessors to select one of the two high-resolution images shown Side-by-Side based on the following evaluation criteria placed in order of their importance:
@@ Appendix D Human evaluation
The assessed high-resolution images are generated from the same low-resolution image. We also provide the initial low-resolution image as a condition to assess potential color distortions. Figure 8 showcases the main labeling interface. Users can also click on any images to open them in full-screen mode. After that, it becomes possible to toggle between two high-resolution images using hotkeys to assess minor differences in an almost pixel-wise manner.
@@
$p\mathrm{-value}=\mathrm{binom\_test}((n_{\mathrm{wins},1}+n_{\mathrm{eq}}//2,n_{\mathrm{wins},2}+n_{\mathrm{eq}}//2),\;p=0.5),$
@@
Appendix E Dataset preparation
We start from a proprietary pool of several billion image-text pairs, initially collected to train image-text models. The data pool originally contained samples with various quality images, texts, and image-text relevance. The following filtering procedure aims to select high-quality samples suitable for training competitive Super Resolution models.
@@ Appendix E Dataset preparation
We use OpenCLIP [11, 2] ViT-G/14 to obtain image-text relevance scores.
@@
Appendix F Super Resolution of synthetic images
To test the robustness of Super Resolution models on the out-of-distribution (OOD) data, we apply both models for upscaling of 4$\times$ downscaled 1024px generations produced by YaART [17] and SDXL [33] models as two examples of modern text-conditional generative models. The former is a cascaded pixel Diffusion model. The latter operates in latent space, and the generated sample is transformed to pixel space with the help of a decoder network.
We generate samples using prompts from DrawBench [40] and YaBasket [17] with both models. (Add details about the generation setup - sampler, number of steps, condition scale to Appendix). LR images are obtained by downscaling 1024px generations to 256px via cv2.INTER_AREA.Even though both models produce photorealistic samples of high quality, generated images could still be different from the data observed during training. The discrepancy between training and evaluation data may negatively affect the performance of SR models and cause artifacts.
@@ Results
Below, we provide only qualitative evaluation since there is no actual ground truth for this setting, as generative models produce outputs that only approximate real-world images. Hence, the calculation of no-reference and full-reference metrics is not very meaningful. Representative examples are shown on Figures 9,10, 11, 12. Models trained without augmentations appeared to be robust enough for evaluation on synthetic images while providing sharper edges and better representing high-frequency details.
@@
Appendix G Acknowledgements
The authors would like to thank Dmitry Baranchuk for the review of the manuscript and thoughtful discussions during the project.

### v2

@@ Appendix B Qualitative Comparison
Appendix C Architecture Details
@@
Appendix D Training Details
We use a linear warmup schedule over 1000 training steps to train the diffusion model on image crops, followed by a constant learning rate of 3e-5. For full-resolution inputs, we decrease the learning rate to 1e-6 and increase the duration of the warmup phase to 10k training steps.
@@ Appendix D Training Details
140k iterations for pretraining and training with the adversarial loss for the GAN-based SR model take approximately three days. Training of diffusion model takes about a month with the same resources.
@@
Appendix E Human Evaluation
Our main tool for training early stopping is side-by-side (SbS) comparisons of Super Resolution results between two subsequent checkpoints of the same model. For that, we use an internal crowdsourcing platform with non-expert assessors. Before labeling, each candidate assessor must pass an exam and achieve at least an 80% accuracy rate among a pre-defined set of 20 assignments. After that, we ask assessors to select one of the two high-resolution images shown Side-by-Side based on the following evaluation criteria placed in order of their importance:
@@ Appendix E Human Evaluation
The assessed high-resolution images are generated from the same low-resolution image. We also provide the initial low-resolution image as a condition to assess potential color distortions. Figure 8 showcases the main labeling interface. Users can also click on any images to open them in full-screen mode and zoom in to better consider the finer details. After that, it becomes possible to toggle between two high-resolution images using hotkeys to assess minor differences in an almost pixel-wise manner.
@@
$p\mathrm{-value}=\mathrm{binom\_test}((n_{\mathrm{wins},1}+n_{\mathrm{eq}}//2,n_{\mathrm{wins},2}+n_{\mathrm{eq}}//2),\;p=0.5),$ (10)
@@
Appendix F Dataset Preparation
We start from a proprietary pool of several billion image-text pairs, initially collected to train image-text models. The data pool originally contained samples with various quality of images, texts, and image-text relevance. The following filtering procedure aims to select high-quality samples suitable for training competitive Super Resolution models.
@@ Appendix F Dataset Preparation
We use OpenCLIP [2] ViT-G/14 to obtain image-text relevance scores.
@@
Appendix G Super Resolution of synthetic images
To test the robustness of Super Resolution models on out-of-distribution (OOD) data, we apply both models to upscaling of 4$\times$ downscaled 1024px images generated by YaART [16] and SDXL [34] models as two examples of modern text-conditional generative models. The former is a cascaded pixel diffusion model, while the latter operates in latent space and transforms the generated sample to pixel space using a decoder network.
We generate samples using prompts from DrawBench [42] and YaBasket [16] with both models. LR images are obtained by downscaling 1024px generations to 256px via cv2.INTER_AREA.Even though both models produce photorealistic samples of high quality, generated images could still be different from the data observed during training. The discrepancy between training and evaluation data may negatively affect the performance of SR models and cause artifacts.
@@ Results
Below, we provide only qualitative evaluation since there is no actual ground truth for this setting, as generative models produce outputs that only approximate real-world images. Hence, the calculation of no-reference and full-reference metrics is not very meaningful. Representative examples are shown in Figures  9,10, 11, 12. Models trained without augmentations appeared to be robust enough for evaluation on synthetic images while providing sharper edges and better representing high-frequency details.
@@
Appendix H Additional Details
H.1 Sampler Hyperparameters Setting for diffusion Model
In our experiments, we evaluate the diffusion model using DPM-Solver++(2M) [30] with 13 sampling steps. This setup demonstrated robust outcomes with a relatively modest number of steps, effectively balancing speed and output quality. To ensure this, we evaluated our diffusion model with different samplers (DDIM [45], UNIPC [60]) and a varying number of sampling steps. The results are presented in Table 9. It is evident from the results that increasing the number of steps beyond those used in our DPM-Solver++ baseline (13 steps) does not enhance performance, whereas reducing the number of steps compromises visual quality.
Table 9: Comparison of different sampler settings to the one used in the experiments (DPM-Solver++, 13). Side-by-side results and basic metrics are introduced. The setting used in this work is optimal in terms of speed and output quality.
Sampler, #steps Wins Loses Ties p-value PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP-IQA $\uparrow$
DPM-Solver++, 13 - - - - 26.66 0.748 0.253 0.719
DPM-Solver++, 6 0.192 0.479 0.329 0.0 27.77 0.793 0.229 0.666
DPM-Solver++, 64 0.287 0.283 0.430 1.0 26.28 0.730 0.275 0.737
UniPC, 6 0.138 0.354 0.508 0.001 27.58 0.786 0.232 0.657
UniPC, 13 0.367 0.338 0.295 0.70 26.50 0.740 0.259 0.728
UniPC, 32 0.267 0.254 0.479 0.85 26.28 0.730 0.274 0.740
UniPC, 64 0.425 0.425 0.150 1.0 26.26 0.730 0.277 0.742
DDIM, 100 0.358 0.354 0.288 1.0 26.60 0.745 0.256 0.730
DDIM, 500 0.417 0.400 0.183 0.85 26.35 0.731 0.269 0.741
DDIM, 1000 0.421 0.383 0.183 0.61 26.32 0.729 0.272 0.735
H.2 Fine-tuning on Full-Resolution Inputs
Our experiments show that fine-tuning on full-resolution inputs brings no benefit compared to training only on crops. Here, we present metrics computed based on the best checkpoints trained on crops only and full-resolution images (Table 10), as well as a side-by-side comparison (Table 11). This can be explained by the fact that both full-resolution and cropped inputs contain the same low-level information, and the global-level semantic information appears to be irrelevant for the SR task.
Table 10: Comparison of only-crops traing vs fine-tuning on full-resolution images based on basic metrics.
Models PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP-IQA $\uparrow$
GAN (crops) 26.00 0.770 0.208 0.806
GAN (full-res) 28.27 0.811 0.182 0.790
Diffusion (crops) 26.66 0.748 0.253 0.719
Diffusion (full-res) 26.83 0.761 0.236 0.715
Table 11: Side-by-side comparison of only-crops training vs fine-tuning on full-resolution images.
Wins Loses Ties p-value
GAN (crops) vs GAN (full-res) 0.425 0.441 0.133 0.847
Diff (crops) vs Diff (full-res) 0.213 0.320 0.467 0.109
We experimented with different learning rates for the full-resolution fine-tune: 1e-4, 1e-5 for GAN, and 3e-5, 1e-6 for diffusion. While training models with the rates used for training on crops (1e-4 for GAN and 3e-5 for diffusion) degrades generation quality (color-shifting and blurred generation), fine-tuning both models even with the smallest rates for 300k iterations leads to no statistical improvements. We provide the results of human evaluations against crop-trained baselines for all mentioned setups in Table 12.
Table 12: Side-by-side comparison of only-crops training vs fine-tuning on full-resolution images with different learning rates.
Wins Loses Ties p-value
GAN (crops) vs GAN (full-res, 1e-5) 0.425 0.441 0.133 0.847
GAN (crops) vs GAN (full-res, 1e-4) 0.658 0.175 0.167 0.000
Diff (crops) vs Diff (full-res, 1e-6) 0.213 0.320 0.467 0.109
Diff (crops) vs Diff (full-res, 3e-5) 0.212 0.155 0.633 0.410
H.3 Early Stopping with Side-by-side Comparison
In our experiments, we adopted a side-by-side comparison to determine the optimal time for early stopping, when further changes in the output of the SR model become imperceptible to human observers (see Appendix E for more detail). However, automated full-reference metrics are more common in the SR literature, so it is interesting to track their evolution during training. We present the measurement of some full-reference metrics in Table 13. One can observe that during pretraining, PSNR and SSIM continue to slightly improve, whereas for GAN training, the metrics stabilize around a specific value after a certain number of iterations. This indicates that above a certain threshold, further improvement in PSNR/SSIM is hardly discernible.
Table 13: Evolution throughout training of PSNR and SSIM.
L1-pretrain GAN Diff
Evaluation Step PNSR $\uparrow$ SSIM $\uparrow$ PNSR $\uparrow$ SSIM $\uparrow$ PNSR $\uparrow$ SSIM $\uparrow$
20k 28.21 0.834 25.84 0.767 20.57 0.659
40k 28.48 0.840 26.01 0.770 20.89 0.656
60k 28.66 0.843 26.01 0.770 21.33 0.663
100k 28.92 0.846 26.00 0.770 22.25 0.675
140k 29.03 0.847 26.00 0.770 22.89 0.681
220k - - - - 23.92 0.701
300k - - - - 25.66 0.733
460k - - - - 26.03 0.741
620k - - - - 26.66 0.748
H.4 Super Resolution Artifacts
We visualize artifacts typical for GAN-based SR and diffusion-based SR in Figure 13 and 14.
Figure 13: GAN’s oversharpening leads to unnatural textures (e.g. fur or grass) and artifacts.
Figure 14: Diffusion is usually worse at high-frequent details (1 row). Additionally, even after big number of training iterations may generate pale, dim images (rows 2-5).
H.5 Efficiency considerations
Model Size
Different SR approaches may vary drastically in terms of the inference pipeline structure and the sizes of the models involved. The superior quality of one method over another could be outweighed in real-world applications by the amount of memory needed to run inference. When considering the size of SR models with diffusion priors (SUPIR, DiffBIR), we include the size of the UNet, text encoders, and VAE decoder.
  • Diff (631 M) = Denoiser (631 M)
  • GAN (631 M) = Generator (614 M) + Discriminator (17 M)
  • SUPIR (17846 M) = SDXL (3469 M) + ControlNet (1332 M) + LLaVA (13045 M)
  • RealESRGAN (21 M) = Generator (17 M) + Discriminator (4 M)
  • DiffBIR (1683 M) = SDv2.1 (1303 M) + Restoration Module (17 M) + ControlNet (363 M)
  • ResShift (174 M) = Denoiser (119 M) + AutoEncoder (55 M)
One can see that different approaches may differ by orders of magnitude in model scale.
Evaluation Settings
We conducted all inference measurements on 1 NVIDIA Tesla A100 GPU with 80GB of VRAM, batch size 1, PyTorch 2.4.0 [35], and CUDA 12.4. We evaluated performance on a $256\text{px}\rightarrow 1024\text{px}$ image super-resolution task.
As mentioned earlier, for our diffusion model, we used 13 steps to generate upscales, while for other diffusion baselines we adopted the default settings from the corresponding papers and repositories: 4 sampling steps for ResShift, 100 sampling steps for SUPIR, and 50 sampling steps for DiffBIR. Note that for these baselines, the reported time spent to produce one image also includes inference of some extra networks such as VAE, LLaVa, BSRNet, etc.
Both our GAN and RealESRGAN models require 1 network evaluation to generate final images.